OpenStudy (anonymous):

Which expression is the simplified form of √k^17

OpenStudy (anonymous):

@amistre64 Help?

OpenStudy (anonymous):

see atttach

OpenStudy (anonymous):

What attach?

OpenStudy (anonymous):

is it \(\sqrt{k^{17}}\)?

OpenStudy (anonymous):

Yes @satellite73

OpenStudy (anonymous):

two goes in to 17 eight times, with a remainder of one, so your answer is \(k^8\sqrt{k}\)

OpenStudy (anonymous):

hope the method is clear, it always works that way

OpenStudy (anonymous):

What Where did 2 Come from?

OpenStudy (anonymous):

square root that is "second root"

OpenStudy (anonymous):

but there is no 2 in the equation?

OpenStudy (anonymous):

if you had the cube root you would divide the exponent by 3

OpenStudy (anonymous):

OH GOTCHA!! THANKS!

OpenStudy (anonymous):

ah that is because when you write a square root you do not write the index, but is it is two in other words \(\sqrt{a}=\sqrt[2]{a}\) but we omit the two

OpenStudy (anonymous):

Gotcha. Thanks!! and one more?

OpenStudy (anonymous):

for other roots you have to write the index cube roots, fourth roots etc yw

OpenStudy (anonymous):

go ahead we can do it

OpenStudy (anonymous):

√72n^13

OpenStudy (anonymous):

same idea as before, but this time we need to recall that \(72=36\times 2\) and therefore \(\sqrt{72}=\sqrt{36\times 2}=\sqrt{36}\times \sqrt{2}=6\sqrt{2}\) you don't need to write the steps, that was my explanation

OpenStudy (anonymous):

How did the thirty six turn into the 6???

OpenStudy (anonymous):

what is the square root of 36?

OpenStudy (anonymous):

and where did n go?

OpenStudy (anonymous):

Ok gotcha on teh square root thing.

OpenStudy (anonymous):

we are not done yet, i am working one piece at a time

OpenStudy (anonymous):

OH ok. :)

OpenStudy (anonymous):

first we find that \(\sqrt{72}=6\sqrt{2}\) then, just like last time, we say 2 goes in to 13 six times, with a remainder of 1, so \(\sqrt{n^{13}}=n^6\sqrt{n}\)

OpenStudy (anonymous):

K gotcha!

OpenStudy (anonymous):

putting them together you get \[\sqrt{72n^{13}}=6n^6\sqrt{2n}\]

OpenStudy (anonymous):

THANK YOU SOOOO MUCH YOU EXPLAINED IT PERFECTLY!!!!!

OpenStudy (anonymous):

yw